i'm having sooo much trouble with this question, i practiced it at least 3 times and i still get it wrong.
Simplify and solve.
3(x+1)-4(y-1)=13
5(x+2)+2(y+3)=0

Answer by checkley71(8403)

3(X+1)-4(Y-1)=13 3X-3-4Y+4=13 3X-4Y=13-4 3X-4Y=9
5(X+2)+2(Y+3)=0 5X+10+2Y+6=0 5X+2Y=-16 NOW MULTIPLY THIS EQUATION BY 2 & ADD:
10X+4Y=-32
3X-4Y=9
----------------------------
13X=-23
X=-23/13 ANSWER.
5(-23/13+2)+2(Y+3=0
5(-23/12+24/12)+2Y+6=0
5(1/12)+2Y+6=0
5/12+2Y+6=0
2Y=-6-5/12
2Y=(-72-5)/12
2Y=-77/12
Y=-77/12*1/2
Y=-77/24 ANSWER.